find the length and width of a rectangular lot with a perimeter of 70 meters if the length is 5 meters more than the width.
Answer by rfer(16322)

P=2L+2W
W=x=15
L=x+5=20
2x+2(x+5)=70
2x+2x+10=70
subtract 10 from both sides
4x+60
divide both sides by 4
x=15